# Volatility Spillovers and Interconnectedness in OPEC Oil Markets: A Network-Based log-ARCH Approach

### v1

@@ 3.1 Interdependence of Oil Price Returns Among OPEC Countries
The correlation matrix in Table 1 shows strong positive comovement among the log returns of oil prices across the six OPEC countries, indicating high market integration and reflecting OPEC’s role in market stabilization. The highest correlations appear among North African countries, likely because of their shared economic structures and similar exposure to global shocks. Gulf countries also exhibit strong interdependence, consistent with their dominant position within OPEC. In contrast, slightly weaker correlations for Iran and the UAE may stem from geopolitical factors or divergent production policies. These findings suggest that while global shocks affect all members similarly, regional asymmetries persist, supporting the use of both constant and dynamic correlation models, such as CCC and DCC-GARCH (Bollerslev, 1990; Engle, 2002), as well as network-based approaches inspired by spatiotemporal ARCH models (see Otto et al., 2024b).
@@
Algeria 1.0000 0.9518 0.9928 0.9946 0.9529 0.9376
Iran 0.9518 1.0000 0.9515 0.9525 0.9804 0.9677
Libya 0.9928 0.9515 1.0000 0.9934 0.9537 0.9350
Nigeria 0.9946 0.9525 0.9934 1.0000 0.9569 0.9356
Saudi Arabia 0.9529 0.9804 0.9537 0.9569 1.0000 0.9769
UAE 0.9376 0.9677 0.9350 0.9356 0.9769 1.0000
@@ 4 Overview of Time-Series Volatility Models
This section reviews the univariate ARCH and GARCH models, including their logarithmic variants (log-ARCH and log-GARCH), which provide the foundation for modeling time-varying volatility in financial time series. We then extend to multivariate specifications—CCC, DCC, and GO-GARCH—that capture static and dynamic correlations and volatility spillovers across assets, and are subsequently used to construct model-based weight matrices for the Network log-ARCH model.
4.1 Introduction to Univariate ARCH and GARCH-type Model
This subsection defines the univariate ARCH and GARCH models employed as building blocks in the CCC, DCC, and GO-GARCH frameworks. We also introduce the univariate log-ARCH process, which later serves as the foundation for the network-based log-volatility models.
4.1.1 Univariate ARCH and GARCH process
For a univariate time series (i.e., $r=1$), the process $Y_{t}$ is given by
$Y_{t}=h_{t}^{1/2}\varepsilon_{t}\,,$ (1)
In an ARCH($p$) model, the volatility dynamics are specified as
$h_{t}=\alpha_{0}+\sum_{i=1}^{p}\alpha_{i}Y_{t-i}^{2},$ (2)
while the GARCH($p,q$) specification is given by
4.1.2 Log-ARCH and log-GARCH models
$\log h_{t}=\alpha_{0}+\sum_{i=1}^{p}\alpha_{i}\log Y_{t-i}^{2}.$ (4)
This can be extended to a log-GARCH model:
$\log h_{t}-\mu_{h}=\phi(\log h_{t-1}-\mu_{h})+u_{t},$ (6)
@@ 4.2 Multivariate GARCH models
While univariate GARCH models capture the conditional volatility of a single financial series, multivariate extensions are required to model the joint dynamics and cross-dependencies across assets. Bollerslev et al. (1988) introduced the VEC-GARCH model, which generalizes the GARCH framework to the multivariate setting by allowing time-varying covariance and correlation structures. Let $\mathbf{Y}_{t}$ be an $r$-dimensional random vector of asset returns. Then:
@@ 4.2.1 Constant and Dynamic Conditional Correlation GARCH Models
Each $h_{it}$ follows a univariate GARCH$(p,q)$ process, as defined in (3). Its general form is:
@@ 4.2.2 The Generalized Orthogonal GARCH Model
where $\mathbf{Z}\in\mathbb{R}^{r\times r}$ is an invertible matrix and $\mathbf{H}_{t}=\text{diag}(h_{1t},\dots,h_{rt})$ is diagonal with each element following a univariate GARCH(1,1)11(1,1)( 1 , 1 ) process:
@@
$\mathbf{R}_{t}=\mathbf{D}_{t}^{-1}\mathbf{\Sigma}_{t}\mathbf{D}_{t}^{-1},\qquad\mathbf{D}_{t}=\text{diag}(\sqrt{h_{1t}},\dots,\sqrt{h_{rt}}).$ (19)
@@ 5.1 Network log-ARCH models
The network log-ARCH model is an extension of spatiotemporal ARCH models (see Otto et al., 2018, 2024a), specifically designed to handle multivariate time series data in the context of network structures (Mattera and Otto, 2024). It captures both temporal and contemporaneous interactions in the volatility process, similar to the ARCH model, but adapted to a network setting. Unlike the previously reviewed univariate and multivariate GARCH models, these network-based approaches allow for volatility spillovers that occur instantaneously across the network in a GARCH-like manner. That is, the conditional variance in (8) also depends on adjacent nodes in $\mathbf{Y}_{t}$ (e.g., $v=t$), instead of strictly conditioning on all $v<t$. Moreover, multivariate GARCH models are typically not suitable for a large number of cross-sections. This is a significant extension, as it implies that the current volatility is directly influenced by the volatilities of the neighboring nodes.
In a network log-ARCH model, the process is observed at the nodes $V$, with the edges $E$ define the dependence structure in the volatility dynamics. This implies that the volatility at a given node is influenced not only by its own past observations but also by the volatility of its adjacent nodes at the same time point. Formally, the model builds upon the framework proposed by Otto et al. (2024a) for dynamic spatiotemporal ARCH models. The observed process $Y_{t}(s_{i})$ at time $t$ and node $v_{i}$ is given by
@@
where $\mathbf{W}$ is the matrix of edge weights (adjacency matrix) defining the degree of volatility spillovers across the network, $\rho$ is a parameter capturing the strength of these contemporaneous network interactions, and $\mathbf{\Gamma}$ is a diagonal matrix of node-specific ARCH effects [^1] Note that this is a specific extension proposed by Mattera and Otto (2024), deviating from the original spatiotemporal log-ARCH model of Otto et al. (2024a).. The term $\boldsymbol{\omega}$ is constant volatility intercepts. The matrix $\mathbf{W}$, analogous to a spatial weight matrix, defines how the volatility spillovers propagate through the network.
@@
where $\boldsymbol{u}_{t}$ are the log-squared errors, and $\boldsymbol{\phi}_{0}$ is a constant vector. This formulation highlights the endogenous nature of the model, as $\boldsymbol{Y}^{*}_{t}=(\ln(Y^{2}_{t}(v_{i})))_{i\in V}$ appears on both sides of the equation, necessitating specialized estimation techniques such as the Generalized Method of Moments (GMM), as proposed by Otto et al. (2024a) or quasi-maximum-likelihood estimators (QMLE) by Otto et al. (2023) to effectively handle the endogeneity issue.
@@ 5.2 Choice of the Network Weight Matrices and Dissimilarity Measures
The choice of weight matrices $\mathbf{W}$ is based on the similarity between the log returns of the six OPEC countries and model-implied correlations, measured using temporal distance metrics. We consider six alternative dissimilarity measures: Euclidean distances, correlation-based distances, Piccolo distances, and distances derived from the CCC-GARCH, DCC-GARCH, and GO-GARCH models. Three of these were adopted from Mattera and Otto (2024), following the time series clustering methodologies outlined in Maharaj et al. (2019).
@@
$\rho_{ij}^{\text{(corr)}}=\frac{\frac{1}{T}\sum_{t=1}^{T}y_{t}^{(i)}y_{t}^{(j)}-\left(\frac{1}{T}\sum_{t=1}^{T}y_{t}^{(i)}\right)\left(\frac{1}{T}\sum_{t=1}^{T}y_{t}^{(j)}\right)}{\sqrt{\left(\frac{1}{T}\sum_{t=1}^{T}(y_{t}^{(i)})^{2}-\left(\frac{1}{T}\sum_{t=1}^{T}y_{t}^{(i)}\right)^{2}\right)\left(\frac{1}{T}\sum_{t=1}^{T}(y_{t}^{(j)})^{2}-\left(\frac{1}{T}\sum_{t=1}^{T}y_{t}^{(j)}\right)^{2}\right)}}.$ (26)
@@
$\bar{\rho}_{ij}^{\text{(DCC)}}=\frac{1}{T}\sum_{t=1}^{T}\rho_{ij,t}^{\text{(DCC)}},\quad d_{ij}^{\text{(DCC)}}=\sqrt{2\left(1-\bar{\rho}_{ij}^{\text{(DCC)}}\right)}.$ (33)
@@
$w_{ij}^{\text{(DCC)}}=\frac{1}{2\left(1-\bar{\rho}_{ij}^{\text{(DCC)}}\right)+1},\quad w_{ii}^{\text{(DCC)}}=0.$ (34)
@@
$\bar{\rho}_{ij}^{\text{(GO)}}=\frac{1}{T}\sum_{t=1}^{T}\rho_{ij,t}^{\text{(GO)}},\quad d_{ij}^{\text{(GO)}}=\sqrt{2(1-\bar{\rho}_{ij}^{\text{(GO)}})}.$ (35)
@@
$W_{ij}^{\text{(GO)}}=\frac{1}{2(1-\bar{\rho}_{ij}^{\text{(GO)}})+1},\quad W_{ii}^{\text{(GO)}}=0.$ (36)
All raw spatial weight matrices—whether based on Euclidean, correlation, or GARCH-derived dissimilarities—are symmetric by construction, satisfying $W_{ij}=W_{ji}$ and $W_{ii}=0$, thereby ensuring well-defined pairwise comparisons. The GO-GARCH-based dissimilarities offer a complementary perspective to CCC and DCC specifications: while the latter directly model pairwise correlations, GO-GARCH captures time-varying dependence through latent factor structures. This factor-based approach enhances estimation stability in high-dimensional settings and yields smoother inter-country correlation dynamics for network construction.
@@ 6 Results and interpretation
This section presents the study’s empirical findings in three stages. First, we report and discuss the estimation results of the univariate and multivariate GARCH models, which form the basis for the construction of the network weight matrices. Second, we analyze the structural patterns revealed by the network plots, providing insight into the interdependencies among oil-exporting countries. Finally, we evaluate the out-of-sample forecasting performance of each modeling approach by comparing their root mean squared forecast errors (RMSFE) and mean absolute forecast errors (MAFE).
@@ 6.1 Model Estimation
As a foundation for the multivariate GARCH models employed in this study, we first estimate univariate GARCH(1,1) models assuming Student’s t-distributed innovations, with estimated degrees of freedom capturing moderate to strong tail heaviness. This distributional choice is motivated by the heavy-tailed nature of the monthly log-returns of oil prices, as evidenced by the initial analysis of the histograms presented in Figure LABEL:fig:hist.
Table 2: Univariate GARCH(1,1) estimates with Student’s t-innovations for monthly log-returns
Parameter Algeria Iran Libya Nigeria S. Arabia UAE
$\mu$ 0.0037 0.0069 0.0038 0.0029 0.0057 0.0055
(0.2781) (0.0333) (0.2842) (0.4446) (0.1383) (0.1114)
$\omega$ 0.0008 0.0014 0.0007 0.0007 0.0009 0.0008
(0.0766) (0.0270) (0.1689) (0.2630) (0.3305) (0.2962)
$\alpha_{1}$ 0.4432 0.5865 0.4496 0.4553 0.5391 0.5155
(0.0000) (0.0000) (0.0000) (0.0000) (0.0000) (0.0000)
$\beta_{1}$ 0.5408 0.4036 0.5494 0.5437 0.4599 0.4835
(0.0000) (0.0001) (0.0000) (0.0000) (0.0040) (0.0026)
$\nu$ 8.3936 5.4360 8.7382 8.8888 5.0903 5.1511
(0.0019) (0.0001) (0.0019) (0.0023) (0.0000) (0.0000)
LogLik 508.95 512.53 502.36 508.90 546.14 575.59
The univariate GARCH(1,1) estimates in Table 2 provide three key insights that motivate the transition to multivariate volatility modeling. First, all countries exhibit persistent volatility dynamics, with $\alpha_{1}+\beta_{1}\approx 1$, indicating long-lasting shock effects and stationary but highly persistent conditional variances. Second, the heterogeneity in the ARCH coefficients ($\alpha_{1}$) suggests differing short-run volatility responses, likely due to structural or informational asymmetries across markets. Third, the high log-likelihood values confirm the model’s adequacy in capturing volatility clustering and time-varying second moments. The estimated degrees of freedom for the Student’s t-distributed innovations further support the presence of heavy-tailed return distributions. Together, these findings justify the transition to multivariate GARCH models on three empirical grounds: (1) significant $\alpha_{1}$ and $\beta_{1}$ values confirm volatility clustering; (2) cross-country variation in parameters challenges the assumption of constant correlations; and (3) stationary yet persistent volatility ensures stable inputs for modeling interdependence. Hence, the univariate results offer both statistical and substantive foundations for estimating the CCC, DCC, and GO-GARCH models to examine cross-market volatility spillovers.
@@
Table 3 reports the CCC-GARCH and DCC-GARCH estimation results for the monthly log returns of the six OPEC countries, assuming a multivariate Student’s t-distribution. The unconditional means are mostly positive but statistically insignificant, indicating no persistent abnormal returns. In the variance equations, both the ARCH ($\alpha$) and GARCH ($\beta$) coefficients are significant at the 1% level, confirming short- and long-term volatility persistence. The condition $\alpha+\beta<1$ holds across countries, ensuring covariance stationarity. The estimated degrees of freedom, ranging between 5 and 9, further support the appropriateness of modeling innovations using a heavy-tailed multivariate Student’s t-distribution. For correlation dynamics, the DCC parameters ($dcca_{1}=0.1016$, $dccb_{1}=0.8207$) are significant and satisfy $dcca_{1}+dccb_{1}<1$, indicating stable and persistent correlations. Compared to CCC-GARCH, the DCC-GARCH model yields higher log-likelihood and lower AIC/BIC values, supporting its superior fit in capturing the time-varying joint volatility structure across the OPEC markets.
@@
Algeria -0.4324 0.3332 0.3640 -0.5228 -0.5299 -0.1250
Iran 0.2453 0.0099 0.6140 -0.4177 0.6103 0.1250
Libya 0.1985 -0.3691 0.6040 0.4155 -0.4553 0.2810
Nigeria -0.5830 0.2563 0.3190 0.5702 0.3451 -0.2210
Saudi Arabia 0.6094 0.6378 0.1110 0.2144 -0.1425 -0.3790
UAE 0.0461 -0.5294 0.1070 -0.0922 -0.0093 -0.8350
@@
$\omega$ 0.0681 0.1155 0.1545 0.1337 0.1434 0.1055
$\alpha_{1}$ 0.3776 0.2857 0.3056 0.8178 0.5916 0.2721
$\beta_{1}$ 0.6214 0.6175 0.5670 0.1812 0.3973 0.6490

### v2

@@ 3.1 Interdependence of Oil Price Returns Among OPEC Countries
The correlation matrix in Table 2 shows strong positive comovement among the log returns of oil prices across the six OPEC countries, indicating high market integration and reflecting OPEC’s role in market stabilization. The highest correlations appear among the African members, particularly between Algeria, Nigeria, and Libya, likely because of their similar exposure to global oil shocks. Gulf countries show a distinct pattern, as Saudi Arabia and the UAE correlate more strongly with each other and with Iran than with the African members, indicating a degree of intra-Gulf cohesion. In contrast, slightly weaker correlations for Iran and the UAE with the North African countries may stem from geopolitical factors or divergent production policies. These findings suggest that while global shocks affect all members similarly, regional asymmetries persist, supporting the use of both constant and dynamic correlation models, such as CCC and DCC-GARCH (Bollerslev, 1990; Engle, 2002), as well as network-based approaches inspired by spatiotemporal ARCH models (see Otto et al., 2024a).
@@
Algeria 1
Iran 0.9513 1
Libya 0.9927 0.9513 1
Nigeria 0.9945 0.9522 0.9932 1
Saudi Arabia 0.9526 0.9806 0.9534 0.9566 1
UAE 0.9378 0.9677 0.9354 0.9360 0.9769 1
@@ 4 Overview of Time-Series Volatility Models
This section reviews the multivariate specifications—CCC, DCC, and GO-GARCH—that capture static and dynamic conditional correlations across assets, and are subsequently used to construct model-based weight matrices for the Network log-ARCH model.
@@ 4.1 Multivariate GARCH models
Multivariate GARCH models are designed to capture the joint dynamics and cross-dependencies across assets. Bollerslev et al. (1988) introduced the VEC-GARCH model, which generalizes the GARCH framework to the multivariate setting by allowing time-varying covariance and correlation structures. Let $\mathbf{Y}_{t}$ be an $r$-dimensional random vector of asset returns. Then:
@@ 4.1.1 Constant and Dynamic Conditional Correlation GARCH Models
Each $h_{it}$ follows a univariate GARCH(p,q)(p,q) process, as defined in (41). Its general form is:
@@ 4.1.2 The Generalized Orthogonal GARCH Model
where $\mathbf{Z}\in\mathbb{R}^{r\times r}$ is an invertible matrix and $\mathbf{H}_{t}=\text{diag}(h_{1t},\dots,h_{rt})$ is diagonal with each element following a univariate GARCH(1,1)(1,1) process:
@@
$\mathbf{R}_{t}=\mathbf{D}_{t}^{-1}\mathbf{\Sigma}_{t}\mathbf{D}_{t}^{-1},\qquad\mathbf{D}_{t}=\left(\mathbf{\Sigma}_{t}\odot\mathbf{I}\right)^{1/2},$ (13)
where $\odot$ denotes the Hadamard product.
@@ 5.1 Network log-ARCH models
The network log-ARCH model is an extension of spatiotemporal ARCH models (see Otto et al., 2018, 2024b), specifically designed to handle multivariate time series data in the context of network structures (Mattera and Otto, 2024). It captures both temporal and contemporaneous interactions in the volatility process, similar to the ARCH model, but adapted to a network setting. Unlike the previously reviewed univariate and multivariate GARCH models, these network-based approaches allow for volatility spillovers that occur instantaneously across the network in a GARCH-like manner. That is, the conditional variance in (2) also depends on adjacent nodes in $\mathbf{Y}_{t}$ (e.g., $v=t$), instead of strictly conditioning on all $v<t$. Moreover, multivariate GARCH models are typically not suitable for a large number of cross-sections. This is a significant extension, as it implies that the current volatility is directly influenced by the volatilities of the neighboring nodes.
In a network log-ARCH model, the process is observed at the nodes $V$, with the edges $E$ define the dependence structure in the volatility dynamics. This implies that the volatility at a given node is influenced not only by its own past observations but also by the volatility of its adjacent nodes at the same time point. Formally, the model builds upon the framework proposed by Otto et al. (2024b) for dynamic spatiotemporal ARCH models. The observed process $Y_{t}(s_{i})$ at time $t$ and node $v_{i}$ is given by
@@
where $\mathbf{W}$ is the matrix of edge weights (adjacency matrix) defining the degree of volatility spillovers across the network, $\rho$ is a parameter capturing the strength of these contemporaneous network interactions, and $\mathbf{\Gamma}$ is a diagonal matrix of node-specific ARCH effects11endnote: 1Note that this is a specific extension proposed by Mattera and Otto (2024), deviating from the original spatiotemporal log-ARCH model of Otto et al. (2024b).. The term $\boldsymbol{\omega}$ is constant volatility intercepts. The matrix $\mathbf{W}$, analogous to a spatial weight matrix, defines how the volatility spillovers propagate through the network.
@@
where $\boldsymbol{u}_{t}$ are the log-squared errors, and $\boldsymbol{\phi}_{0}$ is a constant vector. This formulation highlights the endogenous nature of the model, as $\boldsymbol{Y}^{*}_{t}=(\ln(Y^{2}_{t}(v_{i})))_{i\in V}$ appears on both sides of the equation, necessitating specialized estimation techniques such as the Generalized Method of Moments (GMM), as proposed by Otto et al. (2024b) or quasi-maximum-likelihood estimators (QMLE) by Otto et al. (2025) to effectively handle the endogeneity issue.
@@ 5.2 Choice of the Network Weight Matrices and Dissimilarity Measures
The weight matrices $\mathbf{W}$ are constructed based on the similarity between the log returns of the six OPEC countries and model-implied conditional correlations, measured using temporal distance metrics. We consider six alternative dissimilarity measures: Euclidean distances, correlation-based distances, Piccolo distances. These three dissimilarity measures are adopted from Mattera and Otto (2024). And we introduce three new conditional correaltion distances derived from the CCC-GARCH, DCC-GARCH, and GO-GARCH models, following the time series clustering methodologies outlined in Maharaj et al. (2019).
@@
$\rho_{ij}^{\text{(corr)}}=\frac{\frac{1}{T}\displaystyle\sum_{t=1}^{T}y_{t}^{(i)}y_{t}^{(j)}-\left(\frac{1}{T}\sum_{t=1}^{T}y_{t}^{(i)}\right)\left(\frac{1}{T}\sum_{t=1}^{T}y_{t}^{(j)}\right)}{\sqrt{\left(\frac{1}{T}\displaystyle\sum_{t=1}^{T}(y_{t}^{(i)})^{2}-\left(\frac{1}{T}\sum_{t=1}^{T}y_{t}^{(i)}\right)^{2}\right)\left(\frac{1}{T}\displaystyle\sum_{t=1}^{T}(y_{t}^{(j)})^{2}-\left(\frac{1}{T}\sum_{t=1}^{T}y_{t}^{(j)}\right)^{2}\right)}}.$ (20)
@@
${\rho}_{ij}^{\text{(DCC)}}=\frac{1}{T}\sum_{t=1}^{T}\rho_{ij,t}^{\text{(DCC)}},\quad d_{ij}^{\text{(DCC)}}=\sqrt{2\left(1-{\rho}_{ij}^{\text{(DCC)}}\right)}.$ (27)
@@
$W_{ij}^{\text{(DCC)}}=\frac{1}{2\left(1-{\rho}_{ij}^{\text{(DCC)}}\right)+1},\quad W_{ii}^{\text{(DCC)}}=0.$ (28)
@@
${\rho}_{ij}^{\text{(GO)}}=\frac{1}{T}\sum_{t=1}^{T}\rho_{ij,t}^{\text{(GO)}},\quad d_{ij}^{\text{(GO)}}=\sqrt{2(1-{\rho}_{ij}^{\text{(GO)}})}.$ (29)
@@
$W_{ij}^{\text{(GO)}}=\frac{1}{2(1-{\rho}_{ij}^{\text{(GO)}})+1},\quad W_{ii}^{\text{(GO)}}=0.$ (30)
All raw spatial weight matrices—whether based on Euclidean, correlation, Piccolo-Based, or MGARCH-derived dissimilarities—are symmetric by construction, satisfying $W_{ij}=W_{ji}$ and $W_{ii}=0$, thereby ensuring well-defined pairwise comparisons. The GO-GARCH-based dissimilarities offer a complementary perspective to CCC and DCC specifications: while the latter directly model pairwise correlations, GO-GARCH captures time-varying dependence through latent factor structures. This factor-based approach enhances estimation stability in high-dimensional settings and yields smoother inter-country correlation dynamics for network construction.
@@ 6 Results and interpretation
This section presents the study’s empirical findings in three stages. First, we report the estimation results of the univariate and multivariate GARCH models using the full sample, which form the basis for constructing the network weight matrices. Second, we analyze the structural patterns revealed by the network plots to provide insight into the interdependencies among oil-exporting countries. Finally, we evaluate out-of-sample forecasting performance using an in-sample training period of 450 out of 503 months. We assess predictive accuracy by comparing Root Mean Squared Forecast Errors (RMSFE) and Mean Absolute Forecast Errors (MAFE), and by employing both standard and heteroscedasticity-robust Diebold-Mariano test statistics. We also apply the Model Confidence Set (MCS) procedure to rank the models from best to worst. The section concludes by discussing the advantages of Network Volatility models over standard MGARCH specifications. Furthermore, a full forecasting sensitivity analysis, assessing model performance across different in-sample training ranges, is provided in Table E.1.
@@ 6.1 Model Estimation
As a foundation for the multivariate GARCH models employed in this study, we first estimate univariate GARCH(1,1) models assuming Student’s t-distributed innovations using the full dataset for this part, with estimated degrees of freedom capturing moderate to strong tail heaviness. This distributional choice is motivated by the non-Gaussian properties of monthly oil price log-returns, as evidenced by the leptokurtic density estimates and heavy-tailed Q-Q plots presented in Appendix Figure B.1 and Figure B.2.
The univariate GARCH(1,1) estimates in Appendix Table C.1 provide three key insights that motivate the transition to multivariate volatility modeling. First, all countries exhibit persistent volatility dynamics, with $\alpha_{1}+\beta_{1}\approx 1$, indicating long-lasting shock effects and stationary but highly persistent conditional variances. Second, the heterogeneity in the ARCH coefficients ($\alpha_{1}$) suggests differing short-run volatility responses, likely due to structural or informational asymmetries across markets. Third, the high log-likelihood values confirm the model’s adequacy in capturing volatility clustering and time-varying second moments. The estimated degrees of freedom for the Student’s t-distributed innovations further support the presence of heavy-tailed return distributions. Together, these findings justify the transition to multivariate GARCH models on three empirical grounds: (1) significant $\alpha_{1}$ and $\beta_{1}$ values confirm volatility clustering; (2) cross-country variation in parameters challenges the assumption of constant correlations; and (3) stationary yet persistent volatility ensures stable inputs for modeling interdependence. Hence, the univariate results offer both statistical and substantive foundations for estimating the CCC, DCC, and GO-GARCH models to examine cross-market volatility spillovers.
@@
The estimation results for the CCC-GARCH and DCC-GARCH models, assuming a multivariate Student’s t-distribution, are presented in Table 3 for the monthly log returns of the six OPEC countries. The unconditional means are mostly positive but statistically insignificant, indicating no persistent abnormal returns. In the variance equations, both the ARCH ($\alpha$) and GARCH ($\beta$) coefficients are significant for most countries, confirming short- and long-term volatility persistence, except for Saudi Arabia and UAE, where only ARCH coefficients are significant. The condition $\alpha+\beta<1$ holds across countries, ensuring covariance stationarity. The estimated degrees of freedom, ranging between 5 and 9, further support the appropriateness of modeling innovations using a heavy-tailed multivariate Student’s t-distribution. For correlation dynamics, the DCC parameters ($dcca_{1}=0.1016$, $dccb_{1}=0.8207$) are significant and satisfy $dcca_{1}+dccb_{1}<1$, indicating stable and persistent correlations. Compared to CCC-GARCH, the DCC-GARCH model yields higher log-likelihood and lower AIC/BIC values, supporting its superior fit in capturing the time-varying joint volatility structure across the OPEC markets.
@@
Algeria -0.4404 0.1440 -0.3351 -0.3607 0.5218 -0.5201
Iran 0.2449 -0.1160 -0.0040 -0.6230 -0.6064 -0.4132
Libya 0.1822 -0.2720 0.3498 -0.6121 0.4591 0.4299
Nigeria 0.5941 -0.2240 0.2464 0.2940 0.3631 -0.5637
Saudi Arabia 0.5994 0.4260 -0.6178 -0.1122 0.1315 0.2183
UAE 0.0239 0.8120 0.5683 -0.0898 0.0063 -0.0925
@@
$\omega$ 0.1233 0.0771 0.1012 0.1516 0.1329 0.1312
$\alpha_{1}$ 0.2882 0.4033 0.2560 0.2796 0.8065 0.5505
$\beta_{1}$ 0.6021 0.5937 0.6651 0.5884 0.1925 0.4316